# Decentralizing Test-time Adaptation under Heterogeneous Data Streams

### v1

@@ Abstract
While Test-Time Adaptation (TTA) has shown promise in addressing distribution shifts between training and testing data, its effectiveness diminishes with heterogeneous data streams due to uniform target estimation. As previous attempts merely stabilize model fine-tuning over time to handle continually changing environments, they fundamentally assume a homogeneous target domain at any moment, leaving the intrinsic real-world data heterogeneity unresolved. This paper delves into TTA under heterogeneous data streams, moving beyond current model-centric limitations. By revisiting TTA from a data-centric perspective, we discover that decomposing samples into Fourier space facilitates an accurate data separation across different frequency levels. Drawing from this insight, we propose a novel Frequency-based Decentralized Adaptation (FreDA) framework, which transitions data from globally heterogeneous to locally homogeneous in Fourier space and employs decentralized adaptation to manage diverse distribution shifts. Interestingly, we devise a novel Fourier-based augmentation strategy to assist in decentralizing adaptation, which individually enhances sample quality for capturing each type of distribution shifts. Extensive experiments across various settings (corrupted, natural, and medical environments) demonstrate the superiority of our proposed framework over the state-of-the-arts.
1 Introduction
@@
Despite their success, the effectiveness of current TTA models is generally constrained within ideal testing conditions – often involving homogeneous test samples with similar types of distribution shifts. While attempts have been made to address dynamic target distributions in continually changing environments [36], they fundamentally presume a uniform target domain at any time point. Their focus remains on enhancing model robustness against regular changes by stabilizing the fine-tuning process either by periodically resetting model weights [27, 28] or by down-weighting samples that deviate from the estimated distribution [26, 17]. Although these model-centric approaches may offer temporary relief, they do not fully recognize the intrinsic heterogeneity of real-world data. In practice, distribution shifts do not necessarily occur gradually over time but can be multifaceted at a single moment, involving heterogeneous and even conflicting shifts that current TTA models fail to adequately capture.
To address this, it is crucial to understand how these heterogeneous distribution shifts impact model adaptation. When a model attempts to adjust simultaneously to multiple diverse and potentially conflicting shifts, it may encounter adaptation conflicts. Specifically, adjustments made to accommodate one type of shift can interfere with adaptations for another, as different shifts may require conflicting changes to the model parameters. For instance, adapting to variations in image brightness might necessitate parameter updates that conflict with those needed for texture changes. Such conflicts prevent models from generalizing effectively across all encountered shifts, leading to irreversible degradation in predictive capabilities.
Recognizing these issues, we argue for shifting from a model-centric to a data-centric approach that proactively addresses distribution diversity in Fourier space. The rationale is that the frequency domain, unlike the common spatial domain, enables a clearer separation of data variations across different frequency levels. For example, high-frequency components are typically associated with fine-grained features like edges and textures, whereas low-frequency components generally relate to overall structural patterns such as shapes and illumination. By decomposing data into these frequency components, we can effectively isolate and manage different types of distribution shifts. Moreover, since the Fourier transform operates directly on the raw input images at the pixel level, it does not depend on pretrained model outputs, avoiding potential uncertainties due to significant distribution shifts. Importantly, this proactive separation allows us to manage distribution diversity prior to adaptation, offering a robust foundation for subsequent model enhancement.
Building on this insight, we propose a Frequency-based Decentralized Adaptation (FreDA) framework. Specifically, we first dynamically partition incoming data in the Fourier domain using high-frequency information. This initial segmentation facilitates the transition from globally heterogeneous to locally homogeneous data subsets before any model adaptation occurs. On this basis, we propose a decentralized learning strategy that allows multiple local models to independently adjust to their specific data segments while periodically exchanging knowledge to form a cohesive global model. This dual approach not only captures the diversity of distribution shifts to reduce potential conflicting adaptations but also leverages periodic communication among local models to enhance the global model’s generalization across multiple shifts. Furthermore, we introduce a Fourier-based augmentation mechanism paired with an entropy-based sampling strategy, which significantly increases both the quantity and quality of samples for each type of shift. This enhancement further improves the model’s robustness and predictive capabilities in dynamic environments. To summarize, the main contributions of this work are three-fold:
We identify that many existing TTA methods are restricted in a model-centric paradigm that overlooks the data heterogeneity inherent in real-world scenarios. This oversight results in ineffective adaptation when facing diverse distribution shifts simultaneously.
We revisit TTA from a data-centric perspective and introduce the FreDA framework. It reinterprets principles from both Fourier space and decentralized learning, leveraging specialized local adaptations to manage heterogeneous distribution shifts at test time.
We conduct extensive evaluations across corrupted, natural, and medical scenarios – demonstrating its consistent superiority under various circumstances.
2 Preliminaries
@@
Test-Time Adaptation under Mixed Distributions. Test-time adaptation (TTA) aims to adjust a model , initially trained on a source dataset , to a target domain without access to source data or target labels. Traditionally, TTA handles covariate shift by assuming while . This challenge intensifies when includes multiple non-i.i.d sub-distributions , complicating the adaptation process:
@@
Fourier Transformation. Analyzing the frequency components of images is essential for understanding their underlying structures, and Fourier transformation plays a central role in this process. For a single-channel image , its Fourier transformation is given by: where and denote the height and width of the image, respectively, and and are the frequency coordinates. The inverse Fourier transformation allows for reconstructing the original image from its frequency spectrum, efficiently computed using the Fast Fourier Transform (FFT). In the frequency domain, images are characterized by amplitude and phase components, derived from the real and imaginary parts of :
where reveals the intensity of the frequency content, e.g., high-frequency amplitudes highlight edges and fine details while low-frequency amplitudes emphasize the overall structure and gradual changes, and encodes the position of these features within the spatial domain.
3 Connections to Previous Studies
3.1 Non-i.i.d. in Test-time Adaptation
The non-i.i.d. problem in Test-Time Adaptation (TTA) challenges the conventional assumption that target batches are independent and identically distributed (i.i.d.), pushing the boundaries of TTA in real-world scenarios. This issue can be decomposed into two distinct challenges:
Dependent Sampling. This problem arises when the sampling within the target stream is dependent at the class level. Existing methods [45, 9, 47, 32, 22, 48] have addressed this by aiming for class-balanced datasets during model updates, mitigating risks associated with class imbalance over time. They typically adjust sample proportions based on pseudo labels or extend data collection periods to reduce dependent sampling. However, unlike these methods that concentrate on mitigating class-level imbalances, our work focuses on enhancing TTA models in the presence of diverse sample styles or mixed distributions. We address data heterogeneity at the sample level, aiming to improve model adaptation capabilities in face of varying distribution shifts that are not captured by class balancing techniques. Notably, although our method is not tailored for class-dependent issues, our experimental results demonstrate that when class-dependent and mixed distributions coexist, our approach still achieves the best performance – showcasing the broad applicability of our model design.
Mixed Distributions. While attempts have been made to address dynamic target distributions in continually changing environments [36, 45, 26, 28, 17], they fundamentally assume a uniform target domain at each time point. Their approach focuses on strengthening model adaptation to constant changes by stabilizing the fine-tuning process, using periodic model parameter resets [27, 28] or down-weighting samples that deviate from model expectation [26, 28, 17]. These model-centric approaches fail to capture the actual data heterogeneity encountered in practice, causing model degradation in real-world deployment. In contrast, our work re-examines this problem from a data-centric perspective. We manage heterogeneous data streams by decomposing samples into the frequency domain, which facilitates an accurate data separation and allows us to address distribution diversity before adaptation occurs. Although a recent work [27] also consider mixed distribution scenarios, their study targets a broader “Dynamic Wild World” topic without delving deeply into this data heterogeneity problem. Conversely, our study focuses on investigating and managing heterogeneous data streams in TTA and proposing tailored strategies to address this problem.
3.2 Multi-Target Domain Adaptation
Test-time Adaptation under mixed distribution resembles the multi-target unsupervised domain adaptation (MT-UDA) setting [8, 15, 19, 6], where multiple domains exist within the target domain. However, TTA introduces complexities that far exceed those in conventional MT-UDA settings, primarily due to: 1) Inaccessible Labeled Source Data – In TTA, the labeled source distribution is not available, making it challenging to leverage source-target dissimilarities directly. 2) Dynamic and Unpredictable Target Streams – TTA operates on a continuous influx of data, potentially incorporating new, unforeseen distributions, rather than a static, fully observable target dataset. This continuous nature of data flow prevents the establishment of a comprehensive understanding of the target distribution. These constraints complicates the formulation of adaptation strategies that depend on discerning the differences between various subdomains within the targe.
3.3 Decentralized learning, federated learning, and distributed learning
This work also intersects with decentralized, federated, and distributed learning due to our approach of splitting data batches into disjoint subsets and applying decentralized model adaptation: 1) Decentralized learning typically focuses on learning from decentralized, non-i.i.d. data [13]. In this work, however, the data is not originally decentralized; all target samples arrive together, while we proactively split them into disjoint subsets, revealing latent non-i.i.d. characteristics and enabling the effective use of decentralized learning techniques. 2) Federated learning considers data privacy and multi-institutional collaborations within decentralized learning [24]. In our case, as target samples are mixed in a batch, data privacy is not a constraint. However, like federated learning, our approach also involves model collaboration where multiple local models periodically share insights to form a cohesive global model. 3) Distributed learning aims to improve training efficiency on large-scale datasets by partitioning data for synchronized training [23]. In contrast, our method operates in a real-time fine-tuning context with limited data at one time, hence scalability is less of a concern.
3.4 Frequency Domain Learning
In recent years, frequency-based techniques have become integral to transfer learning strategies. For domain generalization, extensive studies apply frequency analysis to gain insights into the learning behavior and model robustness of DNNs [35, 40, 41, 44]. Some approaches leverage data augmentation to improve model performance. They implement adversarial training to identify the most impactful augmentations [18, 21] or introducing generated Fourier-basis functions as additive noise [30, 33] complementary to visual augmentations. In domain adaptation, researchers implement data augmentation processes involving linear interpolation between the amplitude spectra of images with different styles [43, 42, 38, 39]. This methodology effectively reduces domain discrepancies and mitigates the risk of over-fitting to low-level statistical details inherent in amplitude information.
Motivated by these advancements, we observe that frequency information is a potent domain characteristic. This realization, combined with the insight that augmentations to the amplitude component can mitigate overfitting, particularly due to the uniform amplitude distributions resulting from prior clustering, forms the core inspiration for our FreDA approach. While our proposed frequency-domain perturbation strategy shares the underlying idea of manipulating the amplitude component with these prior works, we distinguish ourselves by introducing subtle perturbations directly to the original amplitude components of carefully selected samples. This strategy is better suited to unsupervised fine-tuning task, where the TTA process is sensitive, and excessive augmentation may cause model degradation.
4 TTA under Mixed Distribution Shifts: A Fourier Perspective
4.1 Motivations
Test-Time Adaptation (TTA) methods have been instrumental in managing domain shifts under a single type of target distribution. However, their effectiveness significantly diminishes under scenarios involving multiple distribution shifts. This is evident as models exhibit a marked decrease in sample class discriminability on the same dataset when exposed to mixed target distributions, as demonstrated by comparing Figure 2(a) and (b). A direct approach to addressing this issue involves segregating samples belonging to different target distributions. However, in real-time applications, the specific target subdomains from which incoming samples originate, or whether they conform to the same distribution shifts, are generally agnostic. Attempting to cluster samples based on features extracted by the model can be misleading, as samples from different distributions may exhibit similarities due to belonging to the same category, resulting in poor separability of different target subdomains as shown in Figure 2(c). Interestingly, when clustering is based directly on the high-frequency information of the samples – without relying on model-derived feature extractors – a significant distinction can be made between samples from different target distributions, as shown in Figure 2(d). This observation is not unexpected, considering high-frequency information typically captures variations in image textures and styles, focusing more on the underlying differences in data distributions. Building on the experimental observations and analysis outlined above, the following section proposes leveraging the frequency domain for enhancing the adaptability of TTA methods in more realistic settings involving mixed distribution shifts.
4.2 Frequency-based Decentralized Adaptation
The previous discussions highlight how heterogeneity within target distribution can hinder model adaptation. This raises a natural question: How can we manage this distributional heterogeneity to achieve better adaptation? As established in our earlier section, effectively distinguishing samples associated with different distribution shifts is vital for successful domain adaptation. Moreover, the similarity in high-frequency information of samples provides a strong indication of whether they belong to the same or different target distributions.
Building upon these findings, we tackle the TTA problem by capitalizing on the high-frequency data components and propose a novel Frequency-based Decentralized Adaptation (FreDA) framework. It employs a data-centric approach to partition target samples into multiple homogeneous subdomains in Fourier space, enabling an accurate model adaptation. This strategy is complemented by a novel frequency-based augmentation technique that enriches each target subdomain with augmented samples, thereby further bolstering model adaptation. The overall pipeline of FreDA is detailed in Aglorithm 1.
4.2.1 Frequency-based Decentralized learning
Insight: Fourier transform offers an effective method to extract different frequency components from images, with high-frequency information particularly useful for capturing fine-grained details such as texture and noise. These details often highlight subtle variations among different distribution shifts. By harnessing high-frequency components from images, we can distinguish samples that lead to different distribution shifts within a TTA setting through a simple clustering technique.
@@ 4.2.1 Frequency-based Decentralized learning
Solution: Based on this intuitive insight, we propose a new module called Frequency-based Decentralized Learning. This module leverages frequency information directly extracted from the pixel space to systematically partition data into multiple homogeneous subsets, enabling multiple local models to specialize in capturing each distribution shift individually (see Figure 1). Concurrently, our method enhances collaborative learning by allowing periodic parameter sharing among these local models, thereby boosting the overall model adaptability to diverse distribution shifts.

### v2

# Un-mixing Test-time Adaptation under Heterogeneous Data Streams

@@ Abstract
Deploying deep models in real-world scenarios remains challenging due to significant performance drops under distribution shifts between training and deployment environments. Test-Time Adaptation (TTA) has recently emerged as a promising solution, enabling on-the-fly model adaptation without access to source data. However, its effectiveness degrades significantly in the presence of complex, mixed distribution shifts – common in practical settings – where multiple latent domains coexist. Adapting under such intrinsic heterogeneity, especially in unlabeled and online conditions, remains an open and underexplored challenge. In this paper, we study TTA under mixed distribution shifts and move beyond conventional homogeneous adaptation paradigms. By revisiting TTA from a frequency-domain perspective, we observe that distribution heterogeneity often manifests in Fourier space – for instance, high-frequency components tend to carry domain-specific variations. This motivates us to perform domain-aware separation using high-frequency texture cues, making diverse shift patterns more tractable. To this end, we propose FreDA, a novel Frequency-based Decentralized Adaptation framework that decomposes globally heterogeneous data into locally homogeneous components in the frequency domain. It further employs decentralized learning and augmentation strategies to robustly adapt under complex, evolving shifts. Extensive experiments across various environments (corrupted, natural, and medical) demonstrate the superiority of our proposed framework over the state-of-the-arts.
Index Terms:
Test-time Adaptation, Transfer LearningI Introduction
@@
Despite their success, current TTA models often are limited to ideal testing conditions, typically involving homogeneous test samples with similar types of distribution shifts. In reality, distribution shifts are mixed, overlapping, and even conflicting [11, 12, 13, 14, 15]. For instance, photo management software handles diverse corruptions, such as noise, blur, compression; medical imaging platforms manage inconsistencies from varied acquisition methods; and autonomous driving systems face fluctuating conditions like lighting, weather, and road types.
Such complexity poses serious challenges for existing methods. While recent efforts have extended TTA to continually changing environments [5], they typically assume a uniform target domain at each time step, as illustrated in the continual setting in Figure 1 (a). Some methods periodically reset the model to its source pre-trained state [7, 9] to mitigate the cumulative effect of sequential distribution shifts on adaptation. Others down-weight outlier target samples [6, 10] to suppress the impact of abrupt or rare shifts. Although these strategies aim to improve robustness under dynamic conditions, they fall short when facing the entangled / mixed real-world distributions (see Figure 1 (b)). Their shortcomings stem from a shared limitation: they lack the capacity to disentangle or localize domain variations. As a result, they often overfit to dominant patterns or suffer from catastrophic forgetting, failing to be deployed in realistic, heterogeneous data streams.
Capturing such intrinsic heterogeneity under unlabeled, online conditions is particularly challenging, and effective adaptation in this setting remains an open problem. To tackle this, we propose shifting from coarse, homogeneous adaptation to a fine-grained strategy that explicitly disentangles heterogeneous shifts. Our approach leverages the frequency information to disentangle domain variations, which naturally separates data features across frequency bands: high-frequency components capture fine details like edges and textures, while low-frequency components represent global structures such as shape and illumination (Section IV). By decomposing inputs accordingly, we better characterize distributional variations and identify diverse shifts. Moreover, the Fourier transform operates directly on raw pixel-level inputs, which shields the method from performance degradation under large domain gaps, and also enhancing adaptability and deployment efficiency in real-world applications.
Building upon this insight, we introduce a framework, termed Frequency-based Decentralized Adaptation (FreDA). FreDA begins by partitioning incoming data in the Fourier domain, where high-frequency components are used to group samples with similar shift characteristics. This transforms globally heterogeneous inputs into locally homogeneous subsets before any adaptation takes place. Based on these partitions, FreDA deploys multiple local models, each assigned to a specific subset. These models adapt independently to their respective data streams while periodically synchronizing through parameter exchange. The aggregated parameters form a shared base model, which is then used to reinitialize the local models, ensuring continual and coordinated adaptation across domains. This approach not only alleviates clustering errors via collaborative knowledge sharing but also allows local models to capture diverse distribution shifts. To further enhance robustness, we introduce a novel Fourier-based augmentation scheme that improves sample quality and strengthens adaptation to shift-specific characteristics.
To summarize, the main contributions of this work are three-fold:
We identify a key limitation of most existing TTA methods – their neglect of real-world data heterogeneity – which leads to suboptimal performance when confronted with mixed and diverse distribution shifts.
We propose FreDA, a frequency-based decentralized adaptation framework that leverages spectral decomposition and localized adaptation to effectively address heterogeneous distribution shifts at test-time.
We validate FreDA through extensive experiments on corrupted, natural, and medical benchmarks, demonstrating consistent improvements over state-of-the-art methods across various TTA scenarios.
II Connections to Previous Studies
II-A Transfer Learning, Domain Adaptation, Test-time Training, and Test-time Adaptation
Deep neural networks often experience performance degradation when deployed in environments different from their training settings, due to distribution shifts between source (training) and target (testing) domains. This issue is extensively studied under the umbrella of transfer learning [16, 17], which aims to transfer knowledge across domains. A key branch of transfer learning is domain adaptation (DA) [18, 19, 20, 21], where models trained on a labeled source domain are adapted to a target domain. Depending on the availability of source and target data, DA can be classified into: Supervised DA using labeled data in both source and target domains during training; Unsupervised DA, which relies on labeled source data and unlabeled target data; and Source-free DA, that adapts to the target domain without access to source data, typically due to privacy or transmission constraints.
| Topic | Source Data | Target Labels | Online Adaptation | Model Agnostic |
| Supervised Domain Adaptation | – | |||
| Unsupervised Domain Adaptation | – | |||
| Source-free Domain Adaptation | – | |||
| Test-time Training | ||||
| Test-time Adaptation |
While traditional DA methods assume access to the whole target distribution during adaptation, such access is not always possible in real-world deployment. To address this, two prominent test-time adaptation paradigms have been proposed: Test-time Training (TTT) [22, 23, 24] introduces an auxiliary task – often self-supervised – during pre-training. At test time, the model optimizes this auxiliary objective to adapt to the target distribution accordingly; Test-time Adaptation (TTA) [3, 7, 8] poses a more demanding and practical scenario, where the model must adapt on-the-fly to the test stream without any prior modification during training. This setting emphasizes the need for rapid, real-time model updates to effectively capture the continuously incoming data. A brief comparison of these adaptation settings is summarized in TABLE I. Our work falls under the TTA setting, where we propose a generalizable and robust adaptation framework that maintains strong performance across diverse distribution shifts and corruptions – without relying on source data or altering the pre-training process.
II-B Non-i.i.d. Test-time Adaptation
While conventional TTA methods primarily focus on single domain shifts under the idealized independent and identically distributed (i.i.d.) assumption, real-world deployment often violates these conditions due to inherent data heterogeneity. This discrepancy has motivated recent efforts to extend TTA to more realistic settings, with two primary challenges:
II-B1 Mixed Domains
Modern TTA methods designed for continual domain shifts [5, 25, 6, 9, 10] have pushed beyond the single adaptation setting. However, these approaches assume that the target domain is uniform at each time point – a special case of mixed domains where the change occurs across distinct time periods. In reality, data streams often involve genuinely mixed domains with heterogeneous distributions present simultaneously. Existing approaches primarily focus on stabilizing model updates through periodic parameter resets [7, 9] or importance weighting of target samples [6, 9, 10], but these techniques are ill-equipped to disentangle and manage the complex interleaving of domain shifts, leading to degraded performance. In contrast, our work adopts a data-centric perspective by explicitly addressing mixed domains via frequency-space decomposition, which facilitates domain separation and proactive distribution alignment prior to model adaptation. While recent efforts like [7] acknowledge mixed distribution challenges under the broader “Dynamic Wild World” paradigm, their unified treatment of diverse real-world factors does not systematically address data heterogeneity. Conversely, our study directly targets the core issue of mixed domains in TTA, introducing tailored strategies for disentangling and adapting to co-occurring heterogeneous distributions.
II-B2 Dependent Sampling
The second challenge stems from class-level dependencies introduced by temporal data sampling. This topic has garnered significant research attention, with a growing body of work [25, 26, 27, 28, 29, 30] actively addressing it through strategies such as pseudo-label-based rebalancing, or extended observation windows. These approaches aim to correct the temporal imbalance caused by skewed class distributions over time. However, unlike this these studies that focus on class-level imbalance, our work diverges by targeting sample-level heterogeneity, which remains underexplored. Specifically, we address style diversity and distributional variation at test-time – a challenge that persist even under class-balanced conditions – to improve model robustness under mixed-domain scenarios.
III Problem Definition of mixed domain shifts
@@
Test-time adaptation (TTA) aims to adjust a model , initially trained on a source dataset , to a target domain on a data stream without accessing source data or target labels. TTA handles covariate shift by assuming while . This challenge intensifies when contains multiple non-i.i.d sub-distributions , such as
@@
| Methods | C10 | C100 | IN |
| TBN (Centralized) | 33.8 | 45.8 | 82.5 |
| TBN (Decentralized) | 28.5 | 43.2 | 77.6 |
IV TTA under Mixed Distribution Shifts: A Fourier Perspective
While Test-Time Adaptation (TTA) methods excel under single-type distribution shifts, their performance degrades catastrophically when facing mixed distribution shifts – a ubiquitous challenge in real-world deployment. As demonstrated on CIFAR-10-C (Figure 1 (b)), leading methods like TENT [3], SAR [7] and RoTTA [25] suffer 16% average accuracy drop compared to single-shift scenarios. Conventional TTA paradigms attempt to mitigate this through passive robustness measures, such as entropy regularization [3], gradient masking [7], or model resetting [9]. These methods treat mixed distribution shifts as irreducible noise, deferring corrective actions until model failures occur – a reactive paradigm that inevitably compromises adaptation efficacy.
To address this, we propose a paradigm shift to proactive distribution management – instead of demanding models to handle arbitrary mixtures, we first resolve the heterogeneity via frequency-space decomposition. Our key insight is that mixed distributions exhibit spectral signatures that can be disentangled before model adaptation begins. As evidenced in Figure 2 (c), high-frequency components naturally separate distinct shifts without relying on error-prone model features (see Figure 2 (b). This theoretically grounded decomposition converts the ill-posed mixed-shift adaptation into multiple well-conditioned homogeneous-shift problems, effectively create an ideal adaptation environment from chaotic real-world streams. Crucially, this decomposition enables a decentralized adaptation where we can deploy domain-specific models tailored to high-frequency-separated clusters. Take TBN [31] as an example: while its centralized variant recalculates batch normalization (BN) statistics globally (treating heterogeneous data as a single distribution), our decentralized adaptation computes localized BN parameters within each spectrally disentangled cluster. As shown in Figu 3, this context-aware estimation preserves semantic discriminability: samples exhibit tighter within-class cohesion and sharper between-class boundaries compared to the blurred separability of centralized adaptation. Building on these insights, the following section details how leveraging the frequency domain enhances TTA methods for realistic scenarios with mixed distribution shifts.
V Frequency-based Decentralized Adaptation
The previous discussion highlights how heterogeneity within target distributions can hinder effective model adaptation. This naturally raises the question: How can we manage such distributional heterogeneity to enable better adaptation? As outlined earlier, distinguishing samples associated with different distribution shifts offers a promising pathway. Building on this insight, we address the TTA problem by leveraging high-frequency components in the data and propose a novel framework termed Frequency-based Decentralized Adaptation (FreDA). It partitions target samples into multiple homogeneous subdomains in the Fourier space, facilitating more accurate model adaptation. This is further enhanced by a frequency-based augmentation strategy that enriches each subdomain with diverse samples, strengthening model robustness.
V-A Frequency-based Decentralized Learning
@@
Fourier transform offers an effective method to extract different frequency components from images, with high-frequency information particularly useful for capturing fine-grained details such as texture and noise. These details often highlight subtle variations among different distribution shifts. Based on this insight, we propose a module termed Frequency-based Decentralized Learning. It leverages frequency information extracted from the pixel space to systematically partition data into multiple homogeneous subsets, enabling multiple local models to specialize in capturing each distribution shift individually. Concurrently, our method enables collaborative learning by allowing periodic parameter sharing among these local models via an enhanced base model, thereby boosting overall adaptability to diverse distribution shifts.
V-A1 Frequency Feature Extraction